The length of ladder is 15m and height of the wall is 9m. How much away is the ladder on the ground from the wall?

Answer by Tanya Yadav(2)

a^2=b^2+c^2
x^2=9^2+15^2
x^2=81+205
x^2=306
x=square root of 306
x=17.49
=17.5m